# End-to-End Learning Framework for Solving Non-Markovian Optimal Control

### v4

@@ 2.1 Fractional-Order Derivative
$\left({}^{GL}D_{a+}^{\alpha}f\right)(x)=\lim_{h\to 0^{+}}\frac{1}{h^{\alpha}}\sum_{k=0}^{\left\lfloor\frac{x-a}{h}\right\rfloor}(-1)^{k}\binom{\alpha}{k}f(x-kh),\$ (1)
@@
[…]
@@ 3.1 Mathematical Foundations for Fractional-Order Learning and Control
$G_{k}$ =$=\begin{cases}I&\text{for }k=0,\\ \sum_{j=0}^{k-1}A_{j}G_{k-1-j}&\text{for }k\geq 1,\end{cases}$ (14)
@@
$A_{j}$ =$=\begin{cases}A-\text{diag}(\alpha_{1},\dots,\alpha_{n})&\text{if}j=0,\\ -D(\alpha,j+1)&\text{if }j\geq 1.\end{cases}$ (15)
@@ 4.1 Performance Evaluation on Synthetic Data
[…]
@@
[…]
@@ Appendix B Related Work
Solving the optimal control problem via LQR involves addressing a potentially large linear system with a block Toeplitz matrix of size $\mathbb{R}^{Tn\times Tn}$. While solutions for Toeplitz and block Toeplitz systems have been extensively studied (Trench, 1986; Kalouptsidis et al., 1984; Chandrasekaran & Sayed, 1998), they still present significant computational challenges. Direct solution methods for Toeplitz systems typically exhibit a computational complexity of $\mathcal{O}(N^{2})$, where $N$ represents the degrees of freedom. However, the high memory requirements of these methods often limit their applicability to large-scale problems. Iterative solvers (Chan & Jin, 2007; Strang, 1986) are more memory-efficient and can achieve a complexity of $\mathcal{O}(N\log^{2}(N))$ for a single linear block Toeplitz system. Advanced approaches, such as global and block variants of the generalized minimal residual (GMRES) method (Saad & Schultz, 1986), can efficiently address sequences of block Toeplitz systems with multiple right-hand sides (Jelich et al., 2021). Machine learning-based methods, such as neural operators (Li et al., 2021; Gupta et al., 2021, 2022; Xiao et al., 2022), offer a promising alternative for accelerating the solution of linear systems. Neural operator-assisted Krylov iterations (Luo et al., 2024), for example, have demonstrated significant computational advantages, achieving up to a $\times5$ speedup in computation time and a 16.1×16.1$\times$16.1 × reduction in the number of iterations. These advancements underscore the potential of machine learning techniques to address computational bottlenecks in optimal control for FOLTI systems.
@@ D.1.2 Estimation
$\text{Trace}(K_{B})$ =$\overset{(1)}{=}\frac{n\sigma_{w}^{2}}{N}\,\text{Trace}\left(\frac{I}{p-m-1}\right)$
@@ E.1.1 System Identification
where $\sigma$ is an activation function (e.g., ReLU) and $1,...,L$ indexes the residual blocks.
@@ E.1.2 Temporal Embedding for $A$ Matrix Evolution
where $\lambda_{i}\in\mathbb{C}$ are eigenvalues encoding system stability. A residual MLP processes the concatenated real/imaginary components $[\mathrm{Re}(\Lambda),\mathrm{Im}(\Lambda)]$ to produce base embeddings $A_{0}$.
@@ E.3.1 Cart Pole System Dynamics
[…]

### v5

@@ 2.1 Fractional-Order Derivative
$\left({}^{GL}D_{a+}^{\alpha}f\right)(x)=\lim_{h\to 0^{+}}\frac{1}{h^{\alpha}}\sum_{k=0}^{\left\lfloor\frac{x-a}{h}\right\rfloor}(-1)^{k}\binom{\alpha}{k}f(x-kh),\\$ (1)
@@
where $x\in\mathbb{R}$, 0<h<ba0<h<b-a, and $\alpha>0$. Here, $\binom{\alpha}{k}$ denotes the generalized binomial coefficient, defined as $\binom{\alpha}{k}=\frac{\Gamma(\alpha+1)}{\Gamma(k+1)\Gamma(\alpha-k+1)},$ where $\Gamma(\cdot)$ denotes the gamma function, which is defined as $\Gamma(z)=\int_{0}^{\infty}t^{z-1}e^{-t}\,dt$ for $z>0$.
@@ 3.1 Mathematical Foundations for Fractional-Order Learning and Control
$G_{k}$ $=\begin{cases}I&\text{for }k=0,\\ \sum_{j=0}^{k-1}A_{j}G_{k-1-j}&\text{for }k\geq 1,\end{cases}$ (14)
@@
$A_{j}$ $=\begin{cases}A-\text{diag}(\alpha_{1},\dots,\alpha_{n})&\text{if }j=0,\\ -D(\alpha,j+1)&\text{if }j\geq 1.\end{cases}$ (15)
@@ 4.1 Performance Evaluation on Synthetic Data
Varying fractional orders. By varying the fractional order $\alpha$, we observe that FOLOC framework effectively captures long-term dependencies. The results of our experiments on commensurate FOLTI systems (see Appendix C) with different fractional orders $\alpha=0.1,\;0.3,\;0.5,\;0.7,\;0.9$ are shown in Table 1. As the fractional order $\alpha$ increases, we observe that the MSE and MAE remain approximately the same, with an overall mean MSE of 0.40% (standard deviation: 0.21%) and an overall mean MAE of 0.95% (standard deviation: 0.55%).
Varying time horizons. We show the results of learning fractional-order dynamics with different time horizons T=8, 16, 32, 64, 128T=8,\;16,\;32,\;64,\;128 in Table 1. As the time horizon increases, FOLOC framework demonstrates improved ability to capture the fractional-order dynamics. Compared with at $T=8$, the MSE and MAE at $T=128$ of FOLOC framework decrease by 97.37% and 92.23% respectively.
Varying state dimensions. We vary the system state dimension n=1, 2, 4, 6, 8n=1,\;2,\;4,\;6,\;8 to evaluate its impact on the learning process (assume $n=m$). Increasing the state dimension significantly raises the learning complexity due to the larger number of variables and the interactions or couplings between the states. Despite this, FOLOC framework demonstrates consistent performance, with an overall mean MSE of 0.45% (standard deviation: 0.15%) and an overall mean MAE of 1.24% (standard deviation: 0.36%). Compared with at $n=2$, the MSE (MAE) of FOLOC framework decreases by 13.01% (18.79%) at $n=6$ and 24.06% (25.45%) at $n=8$. This experiment demonstrates the scalability of our model, and verifies its ability to capture the interactions within the fractional dynamics of the state.
@@
Varying training samples. We show the model performance with different number of training samples N=1000, 2000, 4000, 6000, 8000N=1000,\;2000,\;4000,\;6000,\;8000. Compared with at $N=8000$, the MSE only decreases by 1.65% at $N=1000$. This demonstrates that FOLOC framework is sample-efficient and capable of learning the underlying fractional-order dynamics with a limited number of samples.
@@ Appendix B Related Work
Solving the optimal control problem via LQR involves addressing a potentially large linear system with a block Toeplitz matrix of size $\mathbb{R}^{Tn\times Tn}$. While solutions for Toeplitz and block Toeplitz systems have been extensively studied (Trench, 1986; Kalouptsidis et al., 1984; Chandrasekaran & Sayed, 1998), they still present significant computational challenges. Direct solution methods for Toeplitz systems typically exhibit a computational complexity of $\mathcal{O}(N^{2})$, where $N$ represents the degrees of freedom. However, the high memory requirements of these methods often limit their applicability to large-scale problems. Iterative solvers (Chan & Jin, 2007; Strang, 1986) are more memory-efficient and can achieve a complexity of $\mathcal{O}(N\log^{2}(N))$ for a single linear block Toeplitz system. Advanced approaches, such as global and block variants of the generalized minimal residual (GMRES) method (Saad & Schultz, 1986), can efficiently address sequences of block Toeplitz systems with multiple right-hand sides (Jelich et al., 2021). Machine learning-based methods, such as neural operators (Li et al., 2021; Gupta et al., 2021, 2022; Xiao et al., 2022), offer a promising alternative for accelerating the solution of linear systems. Neural operator-assisted Krylov iterations (Luo et al., 2024), for example, have demonstrated significant computational advantages, achieving up to a $5.5\times$ speedup in computation time and a $16.1\times$ reduction in the number of iterations. These advancements underscore the potential of machine learning techniques to address computational bottlenecks in optimal control for FOLTI systems.
@@ D.1.2 Estimation
$\text{Trace}(K_{B})$ $\overset{(1)}{=}\frac{n\sigma_{w}^{2}}{N}\,\text{Trace}\left(\frac{I}{p-m-1}\right)$
@@ E.1.1 System Identification
where $\sigma$ is an activation function (e.g., ReLU) and 1,,L1,...,L indexes the residual blocks.
@@ E.1.2 Temporal Embedding for $A$ Matrix Evolution
where $\lambda_{i}\in\mathbb{C}$ are eigenvalues encoding system stability. A residual MLP processes the concatenated real/imaginary components $\mathrm{Re}(\Lambda),\mathrm{Im}(\Lambda)]$ to produce base embeddings $A_{0}$.
@@ E.3.1 Cart Pole System Dynamics
$\ddot{\theta}$ $=\frac{g\sin\theta+\cos\theta\left(\frac{-m_{p}l\dot{\theta}^{2}\sin\theta-F}{m_{c}+m_{p}}\right)}{l\left(\frac{4}{3}-\frac{m_{p}\cos^{2}\theta}{m_{c}+m_{p}}\right)},$ (116)
where $\ddot{x}$ is the horizontal acceleration of the cart, $\ddot{\theta}$ is the acceleration of the pole, $m_{c}$ is the mass of the cart, $m_{p}$ is the mass of the pole, and $g$ is the acceleration due to gravity. The control input is the external force $F$ applied to the center of mass of the cart.